# Malsight: Exploring Malicious Source Code and Benign Pseudocode for Iterative Binary Malware Summarization

### v1

@@ 6.5.2 Performance From Different Evaluators
After three sets of scores, we obtained the human assessment scores shown in Table VII. Experienced evaluators tend to give more conservative and stable scores, while less experienced evaluators give a wider range of scores with higher variance. Further, we take the arithmetic average of the scores received by these evaluators as the final human assessment score.
TABLE VII: Evaluation of Different Categories of Evaluators
@@
Table b shows the mean value of BLEURT-sum and the mean value of human evaluation on 79 labeled functions. The human evaluation score of 59.87 means that most functions have exceeded 0.5, i.e. the partially usable standard.
TABLE VIII: Real Malware Performance
BLEURT-sum Human Assessment Score AVG time (file)
@@
Figure 10: Real-world Performance. While the ratio of human assessment score to BLEURT-sum remained positive, Malsight is able to obtain higher assessment results.
Further, we formed a data point by combining a BLEURT-sum score with a human evaluation score for the same function’s code summary, thus, we plotted a distribution of 79 data points in Figure 10. It indicates that BLEURT-sum has a positive linear correlation with human evaluation metrics, which shows the rationality of BLEURT-sum. Furthermore, the data distribution indicates that, across the two distinct evaluation metrics, the majority of the evaluation scores for Malsight’s outputs are concentrated within the middle to high range, which shows that our scheme is also usable in the real world.
7 Case Study
In section RQ4, we evaluate Malsight’s performance on 10 real-world malware. To provide more direct evidence and analyze the reasons why Malsight behaves differently on different functions, three functions are presented in this section to discuss the process that Malsight goes through behind the output that is usable, the output that is partially usable, and the output that is unusable.
Figure 11: Redacted pseudocode in Case 1.
@@ 7 Case Study
Malsight’s summary of the function shown in Figure 11 is considered usable. In this case, Malsight summary accurately identifies “setting privileges” as the key point from a lengthy function, which aligns with the task of “privilege escalation” in the human summary.
Figure 12: Redacted pseudocode in Case 2.
@@
Malsight’s summary of the function shown in Figure 12 is considered partially usable. Malsight summary identifies that the function uses an HTTP connection to a specific server and operates on the data, which is consistent with the human summary. However, due to the lack of some background information, it fails to recognize that the specific server is malicious.
Figure 13: Redacted pseudocode in Case 3.
Case 3: TimerFunc.
  • Malsight Summary: Get the command line, the code defines a function that handles or processes data related to a hexadecimal address, retrieves information about Redis commands and modules, Call the NtTib. ExceptionList, the code defines a function that handles or processes data related to a hexadecimal address, retrieves information about Copy the setup.exe, the code defines a function that handles or processes data related to a hexadecimal address, retrieves information about Redis commands and modules Setup the setup, the code defines a function that handles or processes data related to a hexadecimal address, retrieves information about Redis commands and modules.
  • Human Summary: Checking if the autorun.ini file exists; if not, create it and configure it to auto-run setup.exe as specified in the configuration file.
Malsight’s summary of the function shown in Figure 13 is considered unusable. The function is very long, contains numerous function calls, and includes many redundant functions that hinder analysis, resulting in a poor Malsight summary.
8 Discussion
8.1 Ethics
The construction of the dataset MalS and MalP complies with GitHub’s open-source license agreements. We only used open-source repositories with GitHub open-source licenses and downloaded these open-source codes within the rate limit set by GitHub to minimize interference with GitHub’s servers to the lowest extent possible. Additionally, We enlisted volunteers to review the MalS and MalP datasets, as well as the 10 malware samples mentioned in RQ4. With the volunteers’ consent, we adopted their review results as the final dataset.
8.2 Limitations
@@
Real-World Malware vs. Malware Function Summaries: In our exploration of real-world malware, the BLEURT-sum evaluation score of the code summary was approximately 30% lower than the experimental score obtained by MalP in Section 6.2.
The reason could be that our code summarization model lacks the ability to summarize longer functions which occasionally appear in real-world malware. Our analysis shows that about 15% of the functions in the malware contain more than 1000 tokens (while the majority of the functions in the training set stay below 300 tokens), which is likely due to different optimization configurations during compilation. These lengthy function bodies introduce a lot of information and noise, which makes it difficult for the model to extract and summarize the critical code fragments.
Therefore, to better align with real malicious code summaries, researchers should consider doing code summarization work in units of code fragments instead of functions. However, identifying and summarizing important, human-interpretable segments in them requires large amounts of labeled data, or compositional functional insights from dynamic debugging, which poses significant challenges.
@@
In our comparison of assembly code and pseudocode representations, we found that assembly code is challenging for models pre-trained on high-level languages to understand [32]. The structural features of assembly code are completely different from those of high-level languages. Fine-tuning a model using pseudocode can confuse the model’s understanding of function-level structural features, leading to unacceptable error output, possibly due to insufficient datasets.
We train an embedding model specifically for assembly language and build a large dataset at the assembly language level for malware. This method has the potential to outperform general-purpose large models, which typically have a poor understanding of assembly code.
9 Conclusion
In this paper, we introduced Malsight, a novel framework for binary malware summarization by exploring malicious source code and benign pseudocode. The proposed Malsight involves two datasets MalS and MalP, an LLM-based summary model, and an evaluation metrics. Experimental results on three datasets show the effectiveness of the proposed framework. Future work includes the application of the proposed framework to more downstream tasks.
@@ Appendix A Analysis of Traditional Similarity Evaluation Metric
The criteria for evaluating sentence similarity should focus on two key aspects: similar sentences should correspond to higher scores, while dissimilar sentences should correspond to lower scores. Below, we introduce some of the shortcomings of BLEU and other traditional metrics (ROUGE, METEOR) in these aspects.
A.1 Limitations of Overlap Based Methods
@@
METEOR designs on the basis of rouge, following equation 20, 21 and 22.
@@
A.2 Limitations of Embedding Based Methods
In recent years, using word embedding-based methods such as word2vec to determine sentence similarity has become quite popular. For instance, the pseudocode for the word2vec process is illustrated in Algorithm 1.
Algorithm 1 word2vec
1:The reference sentence $ref$ and the candidate sentence $cad$
2:the similarity $score$ of $ref$ and $cad$
3:function Main($cad,ref$)
4:     cad.vector = getVector($cad$)
5:     ref.vector = getVector($ref$)
6:     $\text{score}=\frac{\text{cad}\cdot\text{ref}}{||\text{cad}||\cdot||\text{ref}||}$
7:     return score
8:end function
9:function getVector($sentence$)
10:     words = tokenize(sentence)
11:     sentence.vector = [0,0,..]
[…]
13:         sentence.vector += word.vector
14:     end for
15:     return sentence.vector
16:end function
When calculating the similarity between two sentences, word2vec adds the vectors of each word based on the tokenized results and uses the summed vector as the sentence representation. Finally, cosine similarity is employed to determine the similarity between the sentences. A direct drawback of this approach is that it loses all word order information and heavily depends on the accuracy of the embedding algorithm.
Appendix B Reverse CFG Sorting Algorithm
Algorithm 2 REsort
1:The CFG graph $G_{M}$ for a malware binary file.
2:The reverse topsort list $L_{G}$ of the function call graph.
3:function REsort($G_{M}$)
4:     for each vertex $v$ in $G_{M}.vertices$ do
5:         if $!v.seen$ then
6:              new $G_{M}^{\prime}$
7:              $G_{M}^{\prime}.vertices\leftarrow$ call Tarjan($v$, $G$)
8:              //Tarjan() traverses all strongly connected components in $G_{M}$ and contracts vertices as $G_{M}^{\prime}.vertices$.
9:         end if
10:     end for
11:     $G_{M}^{\prime}\leftarrow$ call BuildTarGraph($G_{M}^{\prime}.vertices$, $G_{M}$)
12:     //BuildTarGraph() build a new directed acyclic graph $G_{M}^{\prime}$ using $G_{M}^{\prime}.vertices$ and $G_{M}.edges$
13:     $L_{G_{M}^{\prime}}\leftarrow$ call RetopSort($G_{M}^{\prime}$)
14:     //RetopSort() obtains the reverse topological sorting sequence of $G_{M}^{\prime}$
15:     $dist[]\leftarrow$ call Dijkstra($G_{M}$)
16:     //Dijkstra() computes the multi-source shortest paths in graph $G_{M}$ from vertices with an in-degree of 0 to each vertix.
17:     $L_{G_{M}}\leftarrow[]$
18:     for each vertex $v^{\prime}$ in $L_{G_{M}^{\prime}}$ do
19:         $mindist\leftarrow\infty$, $idx\leftarrow-1$
20:         //$idx$ determines the starting vertex for DFS in each strongly connected component in $G_{M}$
21:         for each $ver$ in $v^{\prime}.subvertex$ do
22:              if $dist[ver]<mindist$ then
23:                  $mindist\leftarrow dist[ver]$, $idx\leftarrow ver$
24:              end if
25:         end for
26:         $L_{G_{M}}$.append(DFS($idx$, $v^{\prime}$))
27:         //DFS() derives a traversal order in each strongly connected component in $G_{M}$ as part of the final reverse topological sorting.
28:     end for
29:end function
TABLE IX: Performance in different training data proportion
@@ Appendix B Reverse CFG Sorting Algorithm
Algorithm 2 presents our function topological sorting algorithm, with the objective of constructing the reverse topological sorting of the function call graph $G_{M}$. Due to the presence of multi-cycles (i.e., strongly connected components) in $G_{M}$, it is not feasible to perform reverse topological sorting directly. To address this, we utilize the Tarjan algorithm to contract $G_{M}$ into a new directed acyclic graph $G_{M}^{\prime}$, where reverse topological sorting can be applied to obtain an initial function traversal list $L_{G_{M}^{\prime}}$, with each node representing a strongly connected component in $G_{M}$.
Further, to achieve topological sorting for $G_{M}$, we employ the DFS algorithm to sort each node within the strongly connected components. Since our aim is to reverse-sort the entire graph $G_{M}$, with functions closer to the start function preferred towards the end of $L_{G_{M}}$, we utilize the Dijkstra algorithm to compute the multi-source shortest paths $dist[]$ from all nodes with zero in-degree to other nodes (referred to as ’multi-source’ because the function call graph $G_{M}$ may not necessarily be connected). Subsequently, by traversing $L_{G_{M}^{\prime}}$ sequentially, for each strongly connected component, the node with the minimum $dist$ value is selected as the initial node $idx$ for DFS. Leveraging the inherent stack property of DFS, $idx$ ensures its placement at the end of $L_{G_{M}}$. Finally, $L_{G_{M}}$ represents the desired function traversal order.
Appendix C Performance in different proportion of two-phase fine-tuning training data size
We sampled a total of 140,000 pieces of data in different proportions from the two datasets (MalS and BenignC) for two phases of fine-tuning. The results of the evaluation of model performance in each training data proportion are shown in the Table IX. In the experimental part of the text, we choose the ratio of 4:3 with the best effect as the actual training set ratio of Malsight.

### v3

@@ VI-E2 Performance From Different Evaluators
After three sets of scores, we obtain the human assessment scores shown in Table VIII. Experienced evaluators tend to give more conservative and stable scores, while less experienced evaluators provide a wider range of scores with higher variance. Further, we take the arithmetic average of the scores from the evaluators as the final human assessment score.
TABLE VIII: Results of Different Evaluators
@@
Table IX shows the mean value of BLEURT-sum and the mean value of human evaluation on 79 labeled functions. The human evaluation score of 59.87 means that most functions have exceeded the partially usable standard.
TABLE IX: Real Malware Performance
BLEURT-sum Human Evaluation Score AVG time (file)
@@
(a) Scatter Plot
(b) Violin Plot
Figure 10: Real-world Performance. While the ratio of human assessment score to BLEURT-sum remained positive, Malsight can obtain higher assessment results.
Further, we form a data point by combining a BLEURT-sum score with a human evaluation score for the same function’s code summary. Thus, we plot a distribution of 79 data points in Fig. 10(a). It indicates that the BLEURT-sum has a positive linear correlation with human evaluation metrics, which shows its rationality. Furthermore, the data distribution in Fig. 10(b) indicates that, across the two distinct evaluation metrics, the majority of the evaluation scores for Malsight’s outputs are concentrated within the middle to high range, which shows that our scheme is also usable in the real world.
VII Case Study
In Section RQ4, we evaluate Malsight’s performance on ten real-world malware samples. To provide deeper insights into Malsight’s varying performance across functions, we present two functions to illustrate the process behind generating a high-quality output and a lower-quality one.
Figure 11: Pseudocode of case 1&2. The two real malware functions produced varying results, with the model’s performance slightly weaker on the longer function due to length limitations.
@@ VII Case Study
Malsight’s summary shown in Fig. 11(right) is considered high-quality. In this case, Malsight summary accurately identifies “setting privileges” as the key point from a lengthy function, which aligns with the task of “privilege escalation” in the human summary.
@@
Malsight’s summary of the function shown in Fig. 11(left) is considered lower-quality. The summary correctly identifies the function’s use of an HTTP connection to a server and data operations, aligning with the human summary. However, it misses that the server is malicious due to lacking background information. Additionally, the model’s output tends to be lengthy and repetitive due to the function’s large code body. We believe this is related to the model’s parameter size and the input length it can process, using a model with a larger parameter count could further improve Malsight’s performance.
VIII Discussion
@@
Real-World Malware vs. Function Summaries: In our evaluation of real-world malware, the BLEURT-sum score for code summaries was approximately 30% lower than the experimental score from MalP in Section VI-B. This discrepancy may arise from the model’s difficulty in summarizing longer functions, which are more common in real malware. Our analysis reveals that 15.8% of the malware functions exceed 1000 tokens, while most training set functions are under 300 tokens. This length introduces excessive information and noise, complicating the extraction of critical code fragments.
Therefore, we investigate code summaries at the code fragment level, discovering that the model offers enhanced insights in this format. We propose adopting a large model agent approach for generating summaries of code fragments, which can then be integrated to form function-level summaries. Furthermore, utilizing additional labelled data from manual dynamic debugging can further enhance the results.
@@
In our comparison of assembly code and pseudocode representations, we find that assembly code is challenging for models pre-trained on high-level languages to understand [32]. The structural features of assembly code are completely different from those of high-level languages. Fine-tuning a model using pseudocode can confuse the model’s understanding of function-level structural features, leading to unacceptable error output, possibly due to insufficient datasets. On the other hand, the presence of lengthy assembly code can significantly increase the inference overhead for code summarization models. Moreover, the substantial differences in assembly code produced by various compilers may introduce data shifts. This issue is mitigated by consistently treating all code as pseudocode during processing.
IX Conclusion
In this paper, we addressed the challenges of binary malware summarization by introducing Malsight, a novel framework that leverages insights from both malicious source code and benign pseudocode. Malsight was designed to overcome the limitations of existing methods in handling entangled logic and stripped semantics commonly found in malware binaries. Our contributions include the development of two key datasets, MalS and MalP, specifically constructed to facilitate malware summarization. We also introduced MalT5, a lightweight LLM-based model tailored for this task, and proposed BLEURT-sum, an innovative evaluation metric that enhances code summarization performance beyond traditional word overlap metrics. Experimental results on both test sets and real-world malware datasets demonstrated the effectiveness of Malsight compared to other state-of-the-art methods. The framework successfully addressed critical issues such as recovering meaningful summaries from complex and stripped pseudocode, underscoring its robustness in handling real-world challenges.
Future research may focus on improving the model’s ability to generate concise summaries for longer code sequences and applying the Malsight framework to more complex downstream tasks. Additionally, extending the approach to further real-world malware scenarios remains a crucial direction for continued development.
@@ Appendix A Analysis of Traditional Similarity Evaluation Metric
METEOR designs on the basis of rouge, following equation 23, 24 and 25.
@@
TABLE X: Performance in different training data proportion
@@
TABLE XI: Summarization examples on TLS-related Go code by various LLMs.
Method (LLM) Summary
BinT5 Go TLS connection handling with runtime checks.
HexT5 TLS handshake timeout configuration.
WizardCoder-15B This Go code sets up TLS handshake timeouts for HTTP servers, and it ensures secure connections by configuring read/write deadlines during the TLS handshake phase, preventing the server from hanging indefinitely. Additionally, it includes runtime checks to ensure compatibility.
Code Llama-7B The code enforces TLS handshake timeouts to prevent hanging connections, setting deadlines during negotiation. It includes runtime checks to ensure compatibility while keeping encrypted chats from dozing off mid-conversation.
CodeT5+ This code sets TLS timeouts for HTTP servers, ensuring security.
deepseek-llm-7b-chat This code formats an integer, performs string operations, initializes TLS connection, handles time, and conditionally closes a connection or returns a Go interface.
Appendix B Performance in different proportion of two-phase fine-tuning training data size
We sampled a total of 140,000 pieces of data in different proportions from the two datasets (MalS and BenignC) for two phases of fine-tuning. The results of the evaluation of model performance in each training data proportion are shown in the Table X. In the experimental part of the text, we choose the ratio of 4:3 with the best effect as the actual training set ratio of Malsight. Therefore, in the subsequent experiments, we adopt this configuration as the default training set composition for Malsight.
Figure 13: Application of MALSIGHT to other binary code summarization tasks.
Appendix C Case Study Supplement
C-A Malsight for other binary code summarization task
We complement this with a case study on a pseudo-code summarization task on binary files written and compiled in Go (Golang) (shown in Fig. 13), aiming to demonstrate the generalizability of Malsight.
Case: Go Binary Psuedo-code.
  • Malsight Summary: Initiate TLS connection, the code formats an integer into a string, performs various string operations including concatenation, initializes a TLS connection, handles time-related operations such as adding time, and conditionally closes the connection or returns a handle based on runtime checks.
C-B Application of other LLMs to the binary summarzation task
We would like to show the application of other large language models to the binary summarization task. We will use the same pseudo-code snippet (Fig. 13) shown in the previous subsection. The results are presented in Table XI.